OpenStudy (anonymous):

can anyone solve it with all the processes please: lim-->2 x−√x+2/√4x+1−3

OpenStudy (anonymous):

I might be able to help you with this.

OpenStudy (anonymous):

It would help if you could draw your equation. Because there are some parenthesis that might be missing.

OpenStudy (anonymous):

You still there?

OpenStudy (anonymous):

use L'Hopital rule...

OpenStudy (calculusfunctions):

\[\lim_{x \rightarrow 2}\frac{ x -\sqrt{x +2} }{ \sqrt{4x +1}-3 }\]Is this how the question supposed to look, @kabulis ?

OpenStudy (anonymous):

yes this is what my question is..

OpenStudy (calculusfunctions):

If it is, rationalize both the numerator and the denominator. Example: Find\[\lim_{x \rightarrow 0}\frac{ \sqrt{x +1}-\sqrt{2x +1} }{ \sqrt{3x +4}-\sqrt{2x +4} }\] \[=\lim_{x \rightarrow 0}\left( \frac{ \sqrt{x +1}-\sqrt{2x +1} }{ \sqrt{3x +4}-\sqrt{2x +4} } \right)\left( \frac{ \sqrt{x +1}+\sqrt{2x +1} }{ \sqrt{3x +4}+\sqrt{2x +4} } \right)\left( \frac{ \sqrt{3x +4}+\sqrt{2x +4} }{ \sqrt{x +1}+\sqrt{2x +1} } \right)\] \[=\lim_{x \rightarrow 0}\left( \frac{ (x +1)-(2x +1) }{ (3x +4)-(2x +4) } \right)\left( \frac{ \sqrt{3x +4}+\sqrt{2x +4} }{ \sqrt{x +1}+\sqrt{2x +1} } \right)\] \[=\lim_{x \rightarrow 0}\left( \frac{ -x }{ x } \right)\left( \frac{ \sqrt{3x +4}+\sqrt{2x +4} }{ \sqrt{x +1}+\sqrt{2x +1} } \right)\] \[=-\lim_{x \rightarrow 0}\frac{ \sqrt{3x +4}+\sqrt{2x +4} }{ \sqrt{x +1}+\sqrt{2x +1} }\] \[=(-1)\left( \frac{ \sqrt{3(0)+4}+\sqrt{2(0)+4} }{ \sqrt{0+1}+\sqrt{2(0)+1} } \right)\] \[=(-1)\left( \frac{ 2+2 }{ 1+1 } \right)\] \[=-2\]